# Testing for dice control at craps

### v2

@@ 1 Introduction
The basic bet at craps is the pass-line bet. The initial roll is called the come-out roll. The bet is won if 7 or 11 appears (a natural) and lost if 2, 3, or 12 appears (a craps number). Any other number (4, 5, 6, 8, 9, or 10) becomes the point. Once a point has been established, subsequent rolls are called point rolls. If the point reappears before a 7 appears, the bet is won. If a 7 appears before the point reappears, the bet is lost. The latter event is called a seven-out. A winning pass-line bet pays 1 to 1. Once a pass-line bet is resolved, the next roll is a new come-out roll, and the process begins again. Under the assumption of fair dice and no control, the expected gain from a one-unit pass-line bet is $-7/495$.
@@ 2 The Smith–Scott model
The controlled dice shooter sets each of the two dice and therefore has six possible dice sets, $AA$, $AB$, $AC$, $BB$, $BC$, and $CC$. The distribution of dice totals is given by convolution, namely
@@
We believe that this model is conceptually flawed by its implicit assumption that the two dice behave independently. When the two dice are set and thrown together, it seems intuitively clear that independence must fail if $10<\theta<$ 1.
@@
$p_{SS}(x):=[\text{UNIF}[S](\cdot)*\text{UNIF}[S](\cdot)](x)=\frac{6-|x-7|}{36},\$
@@
When the point is 5 or 9, the probability of making the point is maximized by each of the dice sets $AB$, $AC$, and $BC$, with the probability increasing nonlinearly from 2/5 ($\theta=0$) to 1/2 ($\theta=1$).
@@ 3 The Wong–Shackleford model
$\sigma_{1}$ is of order 4, while $\sigma_{2}$ and $\sigma_{3}$ are of order 2. The group $G$ generated by $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ defines a group action on $D$ whose orbits partition $D$ into equivalence classes. As Wong (2005, p. 154) showed, there are 45 equivalence classes, 18 of size 8 and 27 of size 16, accounting for all $6\cdot 4\cdot 6\cdot 4=576$ elements of $D$. The Hard Ways Set #1 and the Sevens Set #1 appear in Wong’s list of the 45 distinct dice sets as $(2,3,2,3)$ and $(1,2,6,5)$ because, to represent a given equivalence class, Wong chose its first element in lexicographic order. The number of equivalence classes, 45, can also be surmised from Burnside’s lemma.
To connect these results with Section 2, we observe that rotation of just the left die, as in
@@
results in an equivalent dice set in the sense of Section 2. The 45 equivalence classes identified above become just six. $AA$, $BB$, and $CC$ are each obtained by combining six equivalence classes of size 8 and one of size 16 (64 elements total), and $AB$, $AC$, and $BC$ are each obtained by combining eight equivalence classes of size 16 (128 elements total). And $3\cdot 64+3\cdot 128=576$.
@@ 4 Reparameterization of the models
We can solve these two equations for $\theta$ in terms of $\eta\in[-7/495,13/40]$ (with $[-7/495,13/40]$ being the intersection of the two $\rho$ ranges) using the quartic and cubic formulas, or using Mathematica. Unfortunately, the resulting formulas for $\theta=\rho^{-1}(\eta)$ are complicated and inelegant, so we do not attempt to reproduce them here, but we will use them in what follows.
@@ 5 Tests based on sample proportions
where $\Phi$ is cumulative distribution function of the standard normal. Its power at $\eta$ is
@@ 6 Tests based on the sample mean $\overline{L}$
[…]
$\bm{P}_{\theta}:=(1-\theta)\frac{1}{36}\begin{pmatrix}12&6&8&10&0\\ 3&27&0&0&6\\ 4&0&26&0&6\\ 5&0&0&25&6\\ 0&0&0&0&36\end{pmatrix}+\theta\frac{1}{16}\begin{pmatrix}4&2&4&6&0\\ 2&12&0&0&2\\ 2&0&12&0&2\\ 3&0&0&11&2\\ 0&0&0&0&16\end{pmatrix},\quad$ (3)
@@
16711968.52551 at $\frac{1671}{196}\approx 8.52551\text{ at }\theta=0\;\;\text{and}\;\;\frac{296}{27}\approx 10.9630\text{ at }\theta=1,$
@@
17687013841646.0407 at $\frac{1768701}{38416}\approx 46.0407\text{ at }\theta=0\;\text{and}\;\frac{63880}{729}\approx 87.6269\text{ at }\theta=1,$
@@
$\bm{P}_{\theta}:=(1-\theta)\frac{1}{36}\begin{pmatrix}12&6&8&10&0\\ 3&27&0&0&6\\ 4&0&26&0&6\\ 5&0&0&25&6\\ 0&0&0&0&36\end{pmatrix}+\theta\frac{1}{36}\begin{pmatrix}14&4&8&10&0\\ 4&30&0&0&2\\ 4&0&30&0&2\\ 6&0&0&28&2\\ 0&0&0&0&36\end{pmatrix}.\quad$ (4)
@@
16711968.52551 at $\frac{1671}{196}\approx 8.52551\text{ at }\theta=0\;\;\text{and}\;\;\frac{306}{13}\approx 23.5385\text{ at }\theta=1,$
@@
17687013841646.0407 at $\frac{1768701}{38416}\approx 46.0407\text{ at }\theta=0\;\text{and}\;\frac{79506}{169}\approx 470.450\text{ at }\theta=1,$
@@
[…]
@@
[…]
@@ 7 The distribution of $L$
where $01>e_{1}(\theta)>e_{2}(\theta)>e_{3}(\theta)>e_{4}(\theta)>$ 0 are the non-unit eigenvalues of $\bm{P}_{\theta}$, and $c_{1}(\theta)>0$, $c_{2}(\theta)<0$, $c_{3}(\theta)<0$, and $c_{4}(\theta)<0$ are suitably defined functions satisfying
@@ 8 The likelihood ratio test
$-2\log\Lambda_{n}(\bm{x})\to_{d}\chi^{2}(1)\text{ as }n\to\infty.$
@@
total =$=\begin{cases}7&\text{if $|\text{pre-total}-7|=5$ and $U_{0}\geq 1-\theta$, or}\\ &\text{if $|\text{pre-total}-7|=3$ and $U_{3}<1/3$ and $U_{0}\geq 1-\theta$},\\ \text{pre-total}&\text{otherwise},\end{cases}$
@@
total =$=\begin{cases}2(\lfloor 4U_{4}\rfloor+2)&\text{if $\text{pre-total}=7$ and $U_{3}<2/3$ and $U_{0}\geq 1-\theta$},\\ \text{pre-total}&\text{otherwise}.\end{cases}$
@@
[…]
0.0 0.0445299 0.0869 0.0441 (±$\pm 0.0021$)
0.025 0.1225830 0.1945 0.1156 (±$\pm 0.0032$)
0.05 0.1998030 0.3556 0.2450 (±$\pm 0.0043$)
0.1 0.3518320 0.7231 0.6118 (±$\pm 0.0049$)
0.2 0.6468240 0.9927 0.9868 (±$\pm 0.0011$)
@@
[…]
0.0 0.0313088 0.1391 0.0733 (±$\pm 0.0026$)
0.025 0.0859974 0.4554 0.3206 (±$\pm 0.0047$)
0.05 0.1398350 0.7987 0.6855 (±$\pm 0.0046$)
0.1 0.2449310 0.9962 0.9932 (±$\pm 0.0008$)
0.2 0.4446420 1.0000 1.0000 (±$\pm 0.0000$)
@@ 9 Conclusions
$01>e_{1}(\theta)>\frac{108}{144}>e_{2}(\theta)>\frac{104+4\,\theta}{144}>e_{3}(\theta)>\frac{100-\theta}{144}>e_{4}(\theta)>$ 0 .
This holds for $10\leq\theta<$ 1, and also for $\theta=1$ except for the third and fourth inequalities, which become equalities. The non-unit eigenvalues are roots of the quartic equation
@@
$01>e_{1}(\theta)>\frac{27+3\,\theta}{36}>e_{2}(\theta)>\frac{26+4\,\theta}{36}>e_{3}(\theta)>\frac{25+3\,\theta}{36}>e_{4}(\theta)>$ 0 .
This holds for $10\leq\theta<$ 1, and also for $\theta=1$ except for the third and fourth inequalities, which become equalities. The non-unit eigenvalues are roots of the quartic equation (13), where
@@
The formulation (15) allows a computer-assisted proof for small and moderate $x$, say $x\leq 75$. The point is that $t(x,\theta)$ is a polynomial in $\theta$ of degree at most $x-1$, hence its derivative with respect to $\theta$ is a polynomial in $\theta$ of degree at most $x-2$. It suffice to show that this derivative is nonnegative. Let us write it as

### v3

@@ 1 Introduction
The basic bet at craps is the pass-line bet. The initial roll is called the come-out roll. The bet is won if 7 or 11 appears (a natural) and lost if 2, 3, or 12 appears (a craps number). Any other number (4, 5, 6, 8, 9, or 10) becomes the point. Once a point has been established, subsequent rolls are called point rolls. If the point reappears before a 7 appears, the bet is won. If a 7 appears before the point reappears, the bet is lost. The latter event is called a seven-out. A winning pass-line bet pays 1 to 1. Once a pass-line bet is resolved, the next roll is a new come-out roll, and the process begins again. Under the assumption of fair dice and no control, the expected gain from a one-unit pass-line bet is 7/495-7/495.
@@ 2 The Smith–Scott model
The controlled dice shooter sets each of the two dice and therefore has six possible dice sets, $AA$, ABAB, ACAC, BBBB, BCBC, and $CC$. The distribution of dice totals is given by convolution, namely
@@
We believe that this model is conceptually flawed by its implicit assumption that the two dice behave independently. When the two dice are set and thrown together, it seems intuitively clear that independence must fail if $0<\theta<1$.
@@
$p_{SS}(x):=[\text{UNIF}[S](\cdot)*\text{UNIF}[S](\cdot)](x)=\frac{6-|x-7|}{36},\\$
@@
When the point is 5 or 9, the probability of making the point is maximized by each of the dice sets $AB$, ACAC, and $BC$, with the probability increasing nonlinearly from 2/5 ($\theta=0$) to 1/2 ($\theta=1$).
@@ 3 The Wong–Shackleford model
$\sigma_{1}$ is of order 4, while $\sigma_{2}$ and $\sigma_{3}$ are of order 2. The group $G$ of order 16 generated by $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{3}$ defines a group action on $D$ whose orbits partition $D$ into equivalence classes. As Wong (2005, p. 154) showed, there are 45 equivalence classes, 18 of size 8 and 27 of size 16, accounting for all $6\cdot 4\cdot 6\cdot 4=576$ elements of $D$. The Hard Ways Set #1 and the Sevens Set #1 appear in Wong’s list of the 45 distinct dice sets as $(2,3,2,3)$ and $(1,2,6,5)$ because, to represent a given equivalence class, Wong chose its first element in lexicographic order. The number of equivalence classes, 45, can also be surmised from Burnside’s lemma:
$|D/G|=\frac{1}{|G|}\sum_{\sigma\in G}|D^{\sigma}|=\frac{(1)576+(6)24+(9)0}{16}=\frac{720}{16}=45.$
(Here $|D/G|$ is the number of orbits and $D^{\sigma}:=\{(a,b,c,d)\in D:\sigma(a,b,c,d)=(a,b,c,d)\}$.)
To connect these results with those of the preceding section, we observe that rotation of just the left die, as in
@@
and reversal of just the left die, as in
$\sigma_{5}(a,b,c,d):=(a,\bar{b},c,d),$
result in an equivalent dice set in the sense of the preceding section. The 45 equivalence classes identified above become just six. $AA$, BBBB, and $CC$ are each obtained by combining six equivalence classes of size 8 and one of size 16 (64 elements total), and $AB$, ACAC, and $BC$ are each obtained by combining eight equivalence classes of size 16 (128 elements total). And Burnside’s lemma gives
$|D/G^{\prime}|=\frac{1}{|G^{\prime}|}\sum_{\sigma\in G^{\prime}}|D^{\sigma}|=\frac{(1)576+(8)24+(119)0}{128}=\frac{768}{128}=6.$
@@ 4 Reparameterization of the models
We can solve these two equations for $\theta$ in terms of $\eta\in[-7/495,13/40]$ (with [7/495,13/40][-7/495,13/40] being the intersection of the two $\rho$ ranges) using the quartic and cubic formulas, or using Mathematica. Unfortunately, the resulting formulas for $\theta=\rho^{-1}(\eta)$ are complicated and inelegant, so we do not attempt to reproduce them here, but we will use them in what follows.
@@ 5 Tests based on sample proportions
where $\Phi$ is the cumulative distribution function of the standard normal. Its power at $\eta$ is
@@ 6 Tests based on the sample mean $\overline{L}$
To see whether a test based on hand-length observations makes sense, we evaluate the mean and variance of the length $L$ of the shooter’s hand, first under the S–S model. These are just the mean and variance of the absorption time of a five-state Markov chain. Its state space is $\{\text{co},\text{p4-10},\text{p5-9},\text{p6-8},\text{7o}\}$, where the states are interpreted as: shooter is coming out, point 4 or 10 is established, point 5 or 9 is established, point 6 or 8 is established, and shooter has sevened out. (We are implicitly using the fact that, regardless of $\theta$, the totals 4 and 10 have equal probabilities; the same is true of 5 and 9 and of 6 and 8.) Its one-step transition matrix is
$\bm{P}_{\theta}:=(1-\theta)\frac{1}{36}\begin{pmatrix}12&6&8&10&0\\ 3&27&0&0&6\\ 4&0&26&0&6\\ 5&0&0&25&6\\ 0&0&0&0&36\end{pmatrix}+\theta\frac{1}{16}\begin{pmatrix}4&2&4&6&0\\ 2&12&0&0&2\\ 2&0&12&0&2\\ 3&0&0&11&2\\ 0&0&0&0&16\end{pmatrix}$$\quad$ (3)
@@
$\frac{1671}{196}\approx 8.52551\text{ at }\theta=0\;\;\text{and}\;\;\frac{296}{27}\approx 10.9630\text{ at }\theta=1,$
@@
$\frac{1768701}{38416}\approx 46.0407\text{ at }\theta=0\;\text{and}\;\frac{63880}{729}\approx 87.6269\text{ at }\theta=1,$
@@
$\bm{P}_{\theta}:=(1-\theta)\frac{1}{36}\begin{pmatrix}12&6&8&10&0\\ 3&27&0&0&6\\ 4&0&26&0&6\\ 5&0&0&25&6\\ 0&0&0&0&36\end{pmatrix}+\theta\frac{1}{36}\begin{pmatrix}14&4&8&10&0\\ 4&30&0&0&2\\ 4&0&30&0&2\\ 6&0&0&28&2\\ 0&0&0&0&36\end{pmatrix}$$\quad$ (4)
@@
$\frac{1671}{196}\approx 8.52551\text{ at }\theta=0\;\;\text{and}\;\;\frac{306}{13}\approx 23.5385\text{ at }\theta=1,$
@@
$\frac{1768701}{38416}\approx 46.0407\text{ at }\theta=0\;\text{and}\;\frac{79506}{169}\approx 470.450\text{ at }\theta=1,$
@@
$\;=\frac{(4-\theta)(8+\theta)(28-\theta)^{2}}{144(10-\theta)(44+\theta)}\quad\text{(S--S model)},$
@@
$\;=\frac{(3-2\,\theta)(196-85\,\theta+6\,\theta^{2})}{27(5-2\,\theta)(11-3\,\theta)}\quad\text{(W--S model)},$
@@ 7 The distribution of $L$
where $1>e_{1}(\theta)>e_{2}(\theta)>e_{3}(\theta)>e_{4}(\theta)>0$ are the non-unit eigenvalues of $\bm{P}_{\theta}$, and $c_{1}(\theta)>0$, $c_{2}(\theta)<0$, $c_{3}(\theta)<0$, and $c_{4}(\theta)<0$ are suitably defined functions satisfying
@@ 8 The likelihood ratio test
$2\log\Lambda_{n}(\bm{x})\to_{d}\chi^{2}(1)\text{ as }n\to\infty.$
@@
[…]
@@
total $=\begin{cases}2(\lfloor 4U_{4}\rfloor+2)&\text{if $\text{pre-total}=7$ and $U_{3}<2/3$ and $U_{0}\geq 1-\theta$},\\ \text{pre-total}&\text{otherwise}.\end{cases}$
@@
$\eta$ $\theta=\rho^{-1}(\eta)0$ power of $\overline{L}$ test of LR test ($\pm\text{s.e.}$)
0.0 0.0445299 0.0869 0.0441 ($\pm 0.0021$)
0.025 0.1225830 0.1945 0.1156 ($\pm 0.0032$)
0.05 0.1998030 0.3556 0.2450 ($\pm 0.0043$)
0.1 0.3518320 0.7231 0.6118 ($\pm 0.0049$)
0.2 0.6468240 0.9927 0.9868 ($\pm 0.0011$)
@@
$\eta$ $\theta=\rho^{-1}(\eta)$ power of $\overline{L}$ test of LR test $(\pm\text{s.e.})$
0.0 0.0313088 0.1391 0.0733 ($\pm 0.0026$)
0.025 0.0859974 0.4554 0.3206 ($\pm 0.0047$)
0.05 0.1398350 0.7987 0.6855 ($\pm 0.0046$)
0.1 0.2449310 0.9962 0.9932 ($\pm 0.0008$)
0.2 0.4446420 1.0000 1.0000 ($\pm 0.0000$)
@@ 9 Conclusions
Finally, for a test of dice control to have credibility, the data collection must be witnessed by independent observers, as was done in the 2004 Wong experiment mentioned in the Introduction. Data could come in the form of a hand-length sample, but more useful would be a sample of dice totals, from which could be inferred the sample mean of hand lengths, the sample proportion of pass-line wins, and the sample proportion of 7s on point rolls. More useful still would be a sample of dice rolls (e.g., instead of recording a 7, one would record $(1,6)$ or $(2,5)$ or $(3,4)$). This would help to determine how well the data fit the model. Needless to say, it would also be useful to know which dice set the shooter is using at each roll.
A significant challenge is obtaining a sufficiently large sample size. Merging samples from multiple experimenters, as was done recently in a large-scale coin-tossing experiment (Bartoš et al., 2025), would be problematic without a consistent dice control technique by the various shooters.
@@
$1>e_{1}(\theta)>\frac{108}{144}>e_{2}(\theta)>\frac{104+4\,\theta}{144}>e_{3}(\theta)>\frac{100-\theta}{144}>e_{4}(\theta)>0.$
This holds for $0\leq\theta<1$, and also for $\theta=1$ except for the third and fourth inequalities, which become equalities. The non-unit eigenvalues are roots of the quartic equation
@@
$1>e_{1}(\theta)>\frac{27+3\,\theta}{36}>e_{2}(\theta)>\frac{26+4\,\theta}{36}>e_{3}(\theta)>\frac{25+3\,\theta}{36}>e_{4}(\theta)>0.$
This holds for $0\leq\theta<1$, and also for $\theta=1$ except for the third and fourth inequalities, which become equalities. The non-unit eigenvalues are roots of the quartic equation (13), where
@@
The formulation (15) allows a computer-assisted proof for small and moderate $x$, say $x\leq 75$. The point is that $t(x,\theta)$ is a polynomial in $\theta$ of degree at most x1x-1, hence its derivative with respect to $\theta$ is a polynomial in $\theta$ of degree at most x2x-2. It suffice to show that this derivative is nonnegative. Let us write it as